OpenStudy (asnaseer):

How to enter text as part of your equation without losing the spaces between the words.

OpenStudy (anonymous):

\text{your text} \[\text{your text}\]

OpenStudy (asnaseer):

Sometimes you may want to have some text as part of the line that contains your equation. If you do this without using the text command, you will end up with something like this:\[x^2+2x+1=0 (some text explanation)\]That was entered as follows:\\([\)x^2+2x+1=0 (some text explanation)\\(]\) To avoid that behaviour, just use the \(\LaTeX\) \(text\) command to give you:\[x^2+2x+1=0\text{ (some text explanation)}\]That was entered as follows:\\([\)x^2+2x+1=0\\(text\){ (some text explanation)}\\(]\) You can also add extra space between the end of the equation and you text by using either the \(quad\) or \(qquad\) \(\LaTeX\) commands to give you something like this:\[x^2+2x+1=0\quad\text{ (some text explanation)}\]\[x^2+2x+1=0\qquad\text{ (some text explanation)}\]Those were entered by have \\(quad\) and \\(qquad\) before the \\(text\) command.

OpenStudy (anonymous):

for single space you may also use just \ for example x\ a \[x\ a\]

OpenStudy (anonymous):

oh nice a latex group where have i been?

OpenStudy (asnaseer):

This group was inspired by myininaya. :)